OpenStudy (anonymous):

How do I simplify ln(1-y^2)=x^2+C. I can solve for see, but I'm trying to figure out how to remove the ln

OpenStudy (owlcoffee):

I'll give you a equality: \[e ^{Lnx}=x\] or, you can apply the definition of logarithm: \[\log_{a} B=C <=> a^C=B\] So applying it: \[\ln(1-y^2)=x^2+c\] \[e ^{x^2+c}=1-y^2\]

OpenStudy (anonymous):

I forgot, the 1-y^2 is under a square root so to remove that would I get 1-y^2=(e^c *e^x^2)^2

OpenStudy (owlcoffee):

You are correct.

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (javalos01):

r u questioning us or is it ur question...just asking

Parth (parthkohli):

\[\ln (x^m) = m \ln(x)\]

OpenStudy (javalos01):

true

Parth (parthkohli):

\[\ln(1 - y^2) = 2x^2 + 2C\]Just a question: does C denote a constant here?

OpenStudy (javalos01):

good job @debins33 n' @ParthKohli

OpenStudy (anonymous):

Yes, c is a constant

Parth (parthkohli):

Then\[\ln(1 - y^2) = 2x^2 + C\]

Parth (parthkohli):

\[1 - y^2 = e^{2x^2 + C} = Ce^{2x^2}\]

OpenStudy (javalos01):

yea its that one

OpenStudy (anonymous):

So then y=sqrt(-Ce^2x^2)+1?

Parth (parthkohli):

Yeah, I guess so, but the + 1 is inside the sqrt.

OpenStudy (anonymous):

Okay, that's right. Thank you

Parth (parthkohli):

Just for the record, I wrote \(e^{2x^2 + C}\) as \(Ce^{2x^2}\) since \(e^C\) is also a constnat.

Parth (parthkohli):

constant* This lag is annoying.

OpenStudy (anonymous):

OKay